Given the thermochemical equations of lead that react to chlorine to produced lead (II) chloride.

Determine the ΔH for 2 PbCl2 → Pb + PbCl4

Pb + Cl2 → PbCl2 ΔH = −223 kJ

PbCl2 + Cl2 → PbCl4 ΔH = −87 kJ


1
Expert's answer
2021-02-28T06:15:41-0500

Solution:

2PbCl2 → Pb + PbCl4, ΔHx = ???

(1st equation): Pb + Cl2 → PbCl2, ΔH1 = −223 kJ

(2nd equation): PbCl2 + Cl2 → PbCl4, ΔH2 = −87 kJ


According to Hess's law, the heat of reaction depends upon Initial and final conditions of reactants and does not depend of the intermediate path of the reaction.


1) Modify the two given equations to get the target equation:

a) (1st eq): flip it so as to put PbCl2 on the reactant side.

b) (2nd eq): do nothing. We need one PbCl4 on the product side and that's what we have.


2) Rewrite the two equations with the changes made (including changes in their enthalpy):

(1*st equation): PbCl2 → Pb + Cl2, ΔH*1 = +223 kJ

(2nd equation): PbCl2 + Cl2 → PbCl4, ΔH2 = −87 kJ


Thus, adding modified equations and canceling out the common species on both sides, we get:

2PbCl2 → Pb + PbCl4


3) Add the ΔH values of (1*st) and (2nd) equations to get your answer:

∆Hx = (+223 kJ/mol) + (-87 kJ/mol) = +136 kJ/mol

Hence, ΔHx is +136 kJ/mol.


Answer: ΔHx is 136 kJ/mol.